# Green-VLA: Staged Vision–Language–Action Model for Generalist Robots Manipulation Team, Sber Robotics Center*

### v1

@@ 3.2.2  Robotics data
Figure 4: Left: Dataset sampling rates used during the R0 phase of GreenVLA training. Right: Number of data samples (frames) per dataset, illustrating relative temporal coverage. The corpus combines large-scale open datasets (e.g., AgibotWorld, DROID, Galaxea) with internally collected humanoid and dexterous-hand data.
@@ 5  Experiment Metrics Across Phases
We evaluate Green-VLA across staged training phases (R0, R1, R2) and heterogeneous embodiments, focusing on task-following, long-horizon robustness, and efficiency. Overall, Green-VLA is a mid-scale $\sim$4B-parameter VLA model. We use PaliGemma (3B) as the vision–language backbone, and couple it with a dedicated flow-matching action expert and lightweight heads that account for the remaining parameters. Our architecture cleanly separates semantic perception from high-frequency control by attaching the action expert on top of the frozen or lightly tuned PaliGemma trunk. For the R0 robotics pretraining stage, we train Green-VLA for over $10^{5}$ optimization steps on a cluster of 48 H100 GPUs, which is sufficient to fully saturate the unified multi-embodiment robotic dataset while remaining substantially lighter in both data and compute than prior very large VLA systems.
@@
For the Simpler benchmarks on WidowX and Google Robot, we compare our R0-pretrained Green-VLA against a range of existing foundation policies, including $\pi_{0}$ (R0 and SFT-fine-tuned), OpenVLA [14], Flower [26], and RT-1X (on RT-1 data). Our R0 model outperforms other methods at the same pure-pretrain stage and achieves performance comparable to several fine-tuned baselines. We run all experiments with the default Simpler horizon (80-200 steps for Google Robot and 80 steps for WidowX) and use our episode-end prediction head (Green-VLA EEP) to terminate execution once the task is judged complete. This is particularly important for Google Robot tasks: unnecessary motion after the goal is reached can easily turn a successful configuration into a failed one, mirroring real-world deployments where the robot should complete the task and then wait for a new instruction instead of “fidgeting” in the scene. For Simpler, we report the mean result aggregated across 7 evaluation.
@@
Model Drawer Move near Pick Coke AVG VM Drawer Move near Pick Coke AVG VA Average
$\pi_{0}$ (Fine-tune) 38.3 65.3 72.7 58.8 25.6 63.7 75.2 54.8 56.8
OpenVLA 35.6 46.2 16.3 27.7 17.7 47.7 54.5 39.8 33.8
RT-1-X 59.7 31.7 56.7 53.4 49.0 32.3 29.7 39.6 46.5
Flower - - - - 42.4
Green-VLA(R0) 62.9 61.2 90.4 71.4 33.5 38.1 75.5 49.1 60.2
Green-VLA(R1) 47.0 58.7 95.0 66.9 34.1 42.9 92.1 56.3 61.3
Green-VLA(R2) 61.0 50.8 98.1 69.9 51.6 71.2 98.2 73.7 71.8
@@
We evaluate DB-MemVLA on WidowX tasks under the same conditions as Green-VLA for 7 runs and aggregate results using default episode lengths of 80 steps.

### v2

@@ 3.2.2  Robotics data
Figure 4: Left: Dataset sampling rates used during the R0 phase of Green-VLA training. Right: Number of data samples (frames) per dataset, illustrating relative temporal coverage. The corpus combines large-scale open datasets (e.g., AgibotWorld, DROID, Galaxea) with internally collected humanoid and dexterous-hand data.
@@ 5  Experiment Metrics Across Phases
We evaluate Green-VLA across staged training phases (R0, R1, R2) and heterogeneous embodiments, focusing on task-following, long-horizon robustness, and efficiency. Overall, Green-VLA is a mid-scale $\sim$5B-parameter VLA model in its latest configuration. We use Qwen3-VL-4B-Instruct as the primary vision–language backbone due to its higher VLM quality, and couple it with a dedicated flow-matching action expert and lightweight heads that account for the remaining parameters. In earlier versions we used PaliGemma (3B), resulting in an overall $\sim$4B-parameter VLA. Our architecture cleanly separates semantic perception from high-frequency control by attaching the action expert on top of the frozen or lightly tuned VLM trunk. For the R0 robotics pretraining stage, we train Green-VLA for over $10^{5}$ optimization steps on a cluster of 64 H100 GPUs, which is sufficient to fully saturate the unified multi-embodiment robotic dataset while remaining substantially lighter in both data and compute than prior very large VLA systems.
@@
For the Simpler benchmarks on WidowX and Google Robot, we compare our R0-pretrained Green-VLA against a range of existing foundation policies, including $\pi_{0}$ (R0 and SFT-fine-tuned), OpenVLA [15], Flower [29], RT-1X [25] (on RT-1 data), X-VLA [45], Magma [42], GR00T-N1.6 [24], DB-MemVLA [12] and EO-1 [26]. Our R0 model outperforms other methods at the same pure-pretrain stage and achieves performance comparable to several fine-tuned baselines. We run all experiments with the default Simpler horizon (80-200 steps for Google Robot and 80 steps for WidowX) and use our episode-end prediction head (Green-VLA EEP) to terminate execution once the task is judged complete. This is particularly important for Google Robot tasks: unnecessary motion after the goal is reached can easily turn a successful configuration into a failed one, mirroring real-world deployments where the robot should complete the task and then wait for a new instruction instead of “fidgeting” in the scene. For Simpler, we report the mean result aggregated across 7 evaluation.
@@
Model Drawer Move near Pick Coke Apple AVG VM Drawer Move near Pick Coke Apple AVG VA Average
$\pi_{0}$ (Fine-tune) 38.3 65.3 72.7 0.0 44.1 25.6 63.7 75.2 0.0 41.1 42.6
$\pi_{0.5}$ (Fine-tune) 57.9 72.5 86.7 0.0 54.3 50.5 73.5 87.4 0.0 52.8 53.6
X-VLA 64.4 84.6 93.7 18.5 65.3 43.7 78.8 96.1 30.7 62.3 63.8
GR00T-N1.6 61.1 73.8 95.3 13.0 60.8 59.5 68.3 89.6 23.3 60.2 60.5
Magma 62.5 68.3 74.3 13.0 54.5 60.3 76.9 70.8 32.8 60.2 57.4
EO-1 71.3 83.8 98.0 52.8 76.5 91.6 81.7 55.0 23.8 63.0 69.8
OpenVLA 35.6 46.2 16.3 0.0 24.5 17.7 47.7 54.5 0.0 30.0 27.2
RT-1-X 59.7 31.7 56.7 40.7 47.2 49.0 32.3 29.7 40.7 37.9 42.6
Flower - - - - - 42.4
Green-VLA    (Paligemma 3B)
Green-VLA(R0) 62.9 61.2 90.4 0.0 53.6 33.5 38.1 75.5 0.0 36.7 45.1
Green-VLA(R1) 47.0 58.7 95.0 0.0 50.1 34.1 42.9 92.1 16.0 46.2 48.1
Green-VLA(R2) 61.0 50.8 98.1 0.0 52.4 51.6 71.2 98.2 28.0 62.3 57.3
Green-VLA    (Qwen3-VL-4B-Instruct)
Green-VLA(R1) 64.8 75.8 85.7 81.5 77.0 35.7 71.9 92.6 66.7 66.7 71.8
@@
Another models we evaluate on WidowX tasks under the same conditions as Green-VLA for 7 runs and aggregate results using default episode lengths of 80 steps.
@@ 7  Contributors and Acknowledgments
Acknowledgments:
We sincerely thank Sber and the Sber Robotics Center for the opportunity to work on such an ambitious project. We are especially grateful to the Body, Brain, Platform, Dev, Managers, and People Support teams for their professionalism, continuous support, and close collaboration—making it possible to bring to life what we once only envisioned.